OpenStudy (anonymous):

The City Housing Authority has received 75 applications from qualified applicants for ten low-income apartments. Five of the apartments are on the north side of town, and five are on the south side. If the apartments are to be assigned by means of a lottery, find the following probabilities. (a) A specific qualified applicant will be selected for one of these apartments. (Round your answer to three decimal places.) (b) Two specific qualified applicants will be selected for apartments on the same side of town. (Round your answer to five decimal places.)

OpenStudy (kropot72):

(a) The required probability is found from: \[P(OneOfAnyOfTen)=\frac {\left(\begin{matrix}1 \\ 1\end{matrix}\right)\left(\begin{matrix}74 \\ 9\end{matrix}\right)}{\left(\begin{matrix}75 \\ 10\end{matrix}\right)}=\frac{74!}{9!65!}\times \frac{10!65!}{75!}=?\]